A monopoly price discriminates between two markets:
Market 1 P1 = 500 - Q1
Market 2 P2 = 300 - Q2
Total Cost = 15000 + 100Q where Q = Q1 + Q2
Required: Determine the quantity and price of each market and the total profit.
1
Expert's answer
2017-11-22T15:06:07-0500
TC=15000+100(Q1+Q2)
MC1=100
MC2=100
TR1=P*Q1=500Q1-Q1^2
TR2=P*Q2=300Q2-Q2^2
MC1=500-2Q1
MC2=300-2Q2
100=500-2Q1
2Q1=400
Q1=200
100=300-2Q2
2Q2=200
Q2=100
P1=500-200=300
P2=300-100=200